# Temporal-Spatial Object Relations Modeling for Vision-and-Language Navigation

### v1

@@ I Introduction
As a crucial step in the advancement of embodied intelligence [1, 2, 3, 4, 5], vision-and-language navigation (VLN) has garnered increasing attention. The goal of VLN [6, 7, 8, 9] is to guide an agent to a designated location, which is described by a natural language instruction. Although related vision-and-language problems have been extensively studied [10, 11, 12, 13], the VLN task remains highly challenging due to the dynamic nature of real-world environments and the complexity of the instructions involved.
@@
Furthermore, we also find that agents often explore some new locations that are not part of the correct path or conduct multiple explorations at the same location. Exploration across new locations is helpful, granting agents critical environmental insights and informing their navigational decisions [19, 33]. However, repetitive revisits to the same viewpoint do not enhance navigational success but rather impair efficiency. To counteract this, we introduce the Turning Back Penalty (TBP) loss function. Specifically, during the training process of the agent, it penalizes the agent each time it passes a previously visited location. This effectively mitigates the issue of revisits, thereby improving navigation efficiency. Our primary contributions can be summarized as follows:
@@ III-B Object Relations
III-B1 Object nouns features.
@@
III-B2 Temporal object relations.
@@
III-B3 Spatial object relations.
@@ III-C Turning back penalty
During the agent’s navigation, it was noted that the agent tended to revisit the same location multiple times. This usually culminates in an elongated navigation path. For instance, an agent’s journey initiates at point $a$ and concludes at point $e$, successfully navigating through the route $a\rightarrow b\rightarrow c\rightarrow d\rightarrow b\rightarrow e$. It can be observed that the direct path composed of $a\rightarrow b\rightarrow e$ would be optimal. The exploration of vertices $c$ and $d$ represents additional exploration by the agent, while revisiting vertex $b$ indicates redundant exploration. Such additional exploration is beneficial as it enables the agent to acquire new knowledge, thereby enhancing its navigational skills. Conversely, redundant exploration does not improve navigation efficiency and leads to unnecessarily prolonged paths, which is disadvantageous.
@@ IV-E Ablation Study
Dataset TOR EOR OSR SR SPL RGS RGSPL
REVERIE $\times$ $\times$ 51.09 46.98 33.75 32.15 23.05
$\times$ 52.85 48.08 33.77 32.55 23.07
$\times$ 55.44 49.22 33.32 33.46 22.90
54.47 49.64 34.56 34.05 23.41
SOON $\times$ $\times$ 50.88 36.17 22.54 6.02 3.75
$\times$ 51.92 38.05 25.31 5.46 3.51
$\times$ 54.28 39.09 25.13 7.08 4.61
56.49 40.71 24.98 7.37 4.56
@@
Dataset TBP OSR SR SPL RGS RGSPL
REVERIE $\times$ 54.47 49.64 34.56 34.05 23.41
55.55 50.30 36.79 35.27 25.98
SOON $\times$ 56.49 40.71 24.98 7.37 4.56
55.46 40.12 26.00 7.67 5.04
@@
In table IV, we observes an intriguing phenomenon. The inclusion of object relationships does indeed significantly enhance the agent’s success rates in navigation (OSR, SR, and RGS). However, there is no pronounced improvement in metrics such as SPL and RGSPL. This suggests that the object relationship module leads the agent to engage in excessive redundant exploration, resulting in elongated navigation paths. Upon integrating the TBP loss function, we observes a significant reduction in the agent’s revisits to the same location. This is illustrated in table V. This change leads to a more efficient task completion and shows clear improvements in metrics such as SPL.
@@ IV-F Qualitative Results
To demonstrate that our method can accurately learn the various relationships between objects, we visualized attention heatmaps of different object interactions. fig. 4(a) showcases the object relationships learned by both the TOR and SOR modules. From this, we observes that both modules are capable of accurately modeling object relationships. The spatial object relations provide a detailed and comprehensive depiction of inherent connections between objects, such as chairs and rooms, or decor and thermostats being closely associated. The spatial object relations supplement the temporal object relations. Certain information not captured by the TOR module is correctly learned in SOR. For instance, the environmental object relationships suggest a connection between decor and bathroom, which aligns with real-world scenarios but is not precisely represented in the trajectory-object relationships.

### v2

@@ I Introduction
In recent years, advancements in vision-and-language navigation (VLN) have demonstrated significant potential to revolutionize intelligent transportation systems, by enabling more intuitive and effective ways for autonomous vehicles to interpret and navigate complex urban environments [1, 2]. The goal of VLN [3, 4, 5, 6] is to guide an agent to a designated location, which is described by a natural language instruction. Although related vision-and-language problems have been extensively studied [7, 8, 9, 10], the VLN task remains highly challenging due to the dynamic nature of real-world environments and the complexity of the instructions involved.
@@
While these modules enhance the agent’s ability to understand and navigate complex environments, they also introduce a new challenge. The detailed tracking and continuous updating of object relations can lead the agent to explore new locations that are not part of the correct path or to conduct multiple explorations at the same location. Exploration across new locations is helpful, granting agents critical environmental insights and informing their navigational decisions [16, 30]. However, repetitive revisits to the same viewpoint do not enhance navigational success but rather impair efficiency. To counteract this, we introduce the Turning Back Penalty (TBP) loss function. Specifically, during the training process of the agent, it penalizes the agent each time it passes a previously visited location. This effectively mitigates the issue of revisits, thereby improving navigation efficiency. Our primary contributions can be summarized as follows:
@@
The rest of the paper is organized as follows. Section II reviews relevant research about VLN. Section III introduces the details of our method. In Section IV, we present the training methodology and parameter settings of our model, and evaluate it on three datasets. We conclude the paper in Section V.
@@ III-B Object Relations
III-B1 Object nouns features
@@
III-B2 Temporal object relations
@@
III-B3 Spatial object relations
@@ III-C Turning back penalty
After introducing the TOR and SOR modules, we observed the agent tended to revisit the same location multiple times. This usually culminates in an elongated navigation path. For instance, an agent’s journey initiates at point $a$ and concludes at point $e$, successfully navigating through the route $a\rightarrow b\rightarrow c\rightarrow d\rightarrow b\rightarrow e$. It can be observed that the direct path composed of $a\rightarrow b\rightarrow e$ would be optimal. The exploration of vertices $c$ and $d$ represents additional exploration by the agent, while revisiting vertex $b$ indicates redundant exploration. Such additional exploration is beneficial as it enables the agent to acquire new knowledge, thereby enhancing its navigational skills. Conversely, redundant exploration does not improve navigation efficiency and leads to unnecessarily prolonged paths, which is disadvantageous.
@@ IV-E Ablation Study
Dataset TOR EOR TL OSR SR SPL RGS RGSPL
REVERIE $\times$ $\times$ 22.12 51.09 46.98 33.75 32.15 23.05
$\times$ 22.31 52.85 48.08 33.77 32.55 23.07
$\times$ 23.07 55.44 49.22 33.32 33.46 22.90
23.52 54.47 49.64 34.56 34.05 23.41
SOON $\times$ $\times$ 36.18 50.88 36.17 22.54 6.02 3.75
$\times$ 37.07 51.92 38.05 25.31 5.46 3.51
$\times$ 36.46 54.28 39.09 25.13 7.08 4.61
39.91 56.49 40.71 24.98 7.37 4.56
@@
Dataset TBP TL OSR SR SPL RGS RGSPL
REVERIE $\times$ 23.52 54.47 49.64 34.56 34.05 23.41
22.00 55.55 50.30 36.79 35.27 25.98
SOON $\times$ 39.91 56.49 40.71 24.98 7.37 4.56
38.75 55.46 40.12 26.00 7.67 5.04
@@
In Table IV, we observes an intriguing phenomenon. The inclusion of object relationships does indeed significantly enhance the agent’s success rates in navigation (OSR, SR, and RGS). However, this enhancement also leads to an increase in the trajectory length (TL) of the agent. As a result, there is no pronounced improvement in metrics such as SPL and RGSPL. This suggests that the object relationship module leads the agent to engage in excessive redundant exploration, resulting in elongated navigation paths. Upon integrating the TBP loss function, we observes a significant reduction in the agent’s revisits to the same location. This is illustrated in Table V. This change leads to a more efficient task completion and shows clear improvements in metrics such as TL, SPL.
@@ IV-F Qualitative Results
To demonstrate that our method can accurately learn the various relationships between objects, we visualized attention heatmaps of different object interactions after training. Fig. 5(a) showcases the object relationships learned by both the TOR and SOR modules. From this, we observes that both modules are capable of accurately modeling object relationships. The spatial object relations provide a detailed and comprehensive depiction of inherent connections between objects, such as chairs and rooms, or decor and thermostats being closely associated. The spatial object relations supplement the temporal object relations. Certain information not captured by the TOR module is correctly learned in SOR. For instance, the environmental object relationships suggest a connection between decor and bathroom, which aligns with real-world scenarios but is not precisely represented in the trajectory-object relationships.